Choose the graph which corresponds to the area above 13.4 for a N(10,2) distribution. -18 2 13.4 22 х
12 134 10 X | 0 10 X 12 13.4
1:14 PM Wed Jul 29 ENO ca 12 13.4 10 х O 어 -26.8 -13.4 13.4 26.8 Х
Answer #1

Answer: Choose the graph which corresponds to the area above 13.4 for a N(10,2) distribution.

Solution:

Z score = x - μ / σ

Z score = 13.4 - 10 / 2

Z score = 1.7

P-value:

P(x >13.4) = 1 - P(x<13.4)

P(x >13.4) = 1 - 0.9554 (from z table)

P(x >13.4) = 0.0446

Therefore,

The second graph is correct answer.